How many moles of aluminum chloride are produced when 92.30 g of aluminum chlorate is heated in a decomposition reaction? 


1
Expert's answer
2021-02-09T03:38:26-0500

The decomposition of aluminum chlorate can be shown as follows:

Al(ClO3)3 = AlCl3 + O2

From here:

m (AlCl3) = m (Al(ClO3)3) × Mr(AlCl3) / Mr(Al(ClO3)3) = 92.30 g × 133.34 g/mol / 277.33 g/mol = 44.38 g


Answer: 44.38 g